OpenStudy (ksaimouli):

(t)(5t^2+1)^-(1/2) drivative

OpenStudy (ksaimouli):

\[(t)(5t^2+1)^\frac{ -3 }{ 2 }\]

random231 (random231):

why dont u use the chain rule!!!

OpenStudy (ksaimouli):

I did, \[(t)(\frac{ -1 }{ 2 })(5t^2+1)^\frac{ -3 }{ 2}(10t)+(5t^2+1)^\frac{ -1 }{ 2 }\]

random231 (random231):

yup its correct so wheres ur problem?

OpenStudy (ksaimouli):

simplify

random231 (random231):

all u can do is take (5t^2+1)^-1/2 common

OpenStudy (ksaimouli):